OpenStudy (anonymous):

clara is climbing up a cliff at of about 8ft per minute what is he rrate in meters

OpenStudy (jdoe0001):

\(\bf \cfrac{8\cancel{ft}}{min}\cdot \cfrac{2.54\cancel{cm}}{\cancel{in}}\cdot \cfrac{12\cancel{in}}{\cancel{ft}}\cdot \cfrac{m}{100\cancel{cm}}\)